OpenStudy (anonymous):

What is the radius of a circle with a sector area of 7 pi sq. ft. and an arc whose measure is 70°? 6 feet 10 feet 18 feet

OpenStudy (e.mccormick):

Do you know the formula for arc area?

OpenStudy (e.mccormick):

If not, it is pretty easy to find it out as a simple percentage of the full circle area.

OpenStudy (e.mccormick):

As you probably know, the area of a cirle is \(\pi r^2\). Well, you can multiply anything by 1 and it is the same thing: \(\pi r^2 \cdot 1\) But what if I make my 1 look like this: \(\pi r^2 \cdot \dfrac{360}{360}\) Still the same thing. BUT there is a new meaning. For 360 degress out of a possible 360 degrees, the area is that. Now, I can find any percentage by replacing the top 360 with the number involved: \(\pi r^2 \cdot \dfrac{70}{360}\) That is the formula for the area of your circle. So set it equal to the area then solve for the radius.